# Daria Sushnikova    George Turkiyyah    Edmond Chow    David Keyes

### v1

@@ 3.4 Complexity analysis
The crucial feature of the $\mathcal{H}^{2}$-MG complexity analysis is the linear complexity of the $\mathcal{H}^{2}$ matrix. According to Remark 1, for the $\mathcal{H}^{2}$ matrix of size $N\times N$, the construction complexity, the memory requirements, and matrix-by-vector product complexity are all $\mathcal{O}(N)$. Assume $c_{H2}$ to be the $\mathcal{H}^{2}$ matrix-by-vector product constant.
First, compute the $\mathcal{H}^{2}$-MG computational complexity $n_{\text{op}}$ to run one V-cycle. The complexity of the fine grid smoothing iterations is $n_{f}c_{H2}N$, the coarse grid smoothing iterations is $n_{c}c_{H2}N_{i}$, for $i=2,\dots,l-1$, the complexity of the direct solver is $c_{d}N_{l}^{3}$, where $c_{d}$ is the direct solver complexity constant, and the restriction and prolongation operator complexity is negligible, compared to the smoothing iterations. The overall complexity is:
@@
Assume, for simplicity, that the block size is $B$ for all blocks and block rank $r$ is fixed for all levels (consider it to be the maximum rank of any block). Also, assume the number of blocks of the initial matrix $A$ is $M$, and the number of blocks stacked together while transferring from level to level is $d$. Thus, $N=MB$, $N_{i}=\frac{Mr}{d^{i-2}}$, for $i=2,\dots,l$, and the overall complexity is:
@@
Using the sum of geometric progression, we obtain:
@@
Now compute the memory requirements. The $\mathcal{H}^{2}$-MG algorithm requires storage of matrices $A$, $A_{i}$ for $i=2,\dots,l$, $U_{i}$, and $V_{i}$ for $i=1,\dots,l-1$. Matrices $U_{i}$ and $V_{i}$ are block-diagonal, and their storage is negligible compared to matrices $A$ and $A_{i}$. Assume $\mathcal{H}^{2}$ memory constant to be $c_{m}$. Thus, the memory requirements are:
@@
Analogously to the time complexity rewriting matrix sizes in terms of $M$, $B$, $r$, and $d$ and applying the sum of geometric progression, we obtain:
@@
Taking into account that $N_{l}$ is a constant, obtain:
@@ 4 Numerical results
The numerical experiments showcase our Python implementation of the $\mathcal{H}^{2}$-MG algorithm, demonstrating $\mathcal{H}^{2}$-MG convergence for two different kernels. The matrices are cast into the $\mathcal{H}^{2}$ format using the MCBH [27, 26] method. The code for the $\mathcal{H}^{2}$-MG algorithm is publicly available at https://github.com/dsushnikova/h2mg.
We compare the $\mathcal{H}^{2}$-MG algorithm with solvers CG and FMM-LU [34]. All three methods benefit from $\mathcal{H}^{2}$ structure of matrix $A$; thus, the comparison is natural. The $\mathcal{H}^{2}$-MG algorithm uses CG iteration for smoothing, illustrating how the coarse-level iterations improve the convergence relative to CG alone. In the plots, we present convergence of the $A$-norm of the error $||e_{k}||_{A}$ on $k$-th iteration. We consider convergence per iteration and per time. Note that the plotted timings include the computation of the $A$-norms. For the CG method we plot all the iterations, for the $\mathcal{H}^{2}$-MG we plot only the fine grid iterations (with the matrix $A$), the coarser levels iterations are seen on the convergence plots as error drops.
In all our tests, we randomly generate $x_{\text{true}}$, then compute $b=Ax_{\text{true}}$ and solve the system $Ax=b$ for $x$.
@@
The right-hand sides were chosen in this manner so that the error norm at each iteration could be easily computed. We also tested a few cases where the right-hand sides were chosen randomly from a standard Gaussian distribution. In these cases, the linear systems were harder to solve, but H2-MG outperformed unpreconditioned CG, just like in the detailed results that we show below.
@@ 4.1.1 Gaussian matrix, analysis of error evolution
In our tests, we randomly generate $x_{\text{true}}$, then compute $b=Ax_{\text{true}}$ and solve the system $Ax=b$ for $x$. During the process, we can compute the error $e_{1*}=x_{k}-x_{\text{true}}$ since we know $x_{\text{true}}$. In the classic multigrid algorithm, the error $e_{1*}$ should behave in a following way: for the fine-grid iterations, the high-frequency components of $e_{1*}$ should decrease rapidly, and during the coarse-grid iterations, successive lower frequency components should decrease in turn.
In this section, we study the behavior of the error $e_{1*}$ in the $\mathcal{H}^{2}$-MG method. Consider $U_{1}$, the transfer matrix from the finest level to the coarser one. $U_{1}$ has orthonormal columns. Assume that the matrix $Q_{1}$ has columns that span the complimentary space.
To analyze the components of the error after smoothing in the basis of the interpolation matrix $U_{1}$ or in other words the part of the error that will be projected to the coarser levels, we compute $\text{pinv}(U_{1})e_{1*}$. Similarly, for the part of the after-smoothing error in the basis of the matrix $Q_{1}$, $\text{pinv}(Q_{1})e_{1e}$.
We study 1D Gaussian matrix with points equally spaced between 00 and $1$, $N=1024$, $\sigma=0.01$, $c=10^{-3}$. The results of this analysis are presented in Figure 5. The blue curves represent the error components after the first smoothing step in the initial $V$-cycle, the red curves are for the second V-cycle.
@@
The graphs clearly demonstrate that the error component $\text{pinv}(U_{1})e_{1*}$ is larger than that in the error component $\text{pinv}(Q_{1})e_{1*}$ components, providing compelling evidence that the CG smoothing can complement coarse grid correction effectively.
4.1.2 Gaussian matrix, analysis of number of smoothing steps
@@
Let us first study the convergence rate of the $\mathcal{H}^{2}$-MG depending on the number of fine iterations. Figure 6 presents the convergence comparison for different numbers of fine iterations, $n_{f}$ to the accuracy $\varepsilon=10^{-9}$, where $\varepsilon=\frac{(eAe^{\top})^{\frac{1}{2}}}{||b||_{2}}<10^{-9}$. The number of coarse iterations is fixed $n_{c}=40$. Fine and coarse iteration comparisons are given for matrix size $N=8\times 10^{4}$.
(a) $A$-norm per iteration
(b) $A$-norm per time
@@
We can see that in terms of convergence per iteration, fewer fine-grid iterations lead to better convergence. Figure 7 presents the convergence comparison to the accuracy $\varepsilon=10^{-9}$ for different numbers of coarse iterations, $n_{c}$. The number of fine iterations is fixed at $n_{f}=1$.
(a) $A$-norm per iteration
(b) $A$-norm per time
@@
4.1.3 Gaussian matrix, asymptotics analysis
In this subsection we consider the asymptotic behavior of the system (9) with matrix $A$ given by (10) for various combinations of parameters $c$ and $\sigma$. Figures 8 and 9 show the convergence of the $\mathcal{H}^{2}$-MG algorithm compared to CG across various problem sizes. The method tolerance is $\varepsilon=10^{-9}$. For the test we use parameters $n_{f}=1$, $n_{c}=40$.
(a) $c=10^{-3}$
(b) $c=10^{-5}$
(a) $c=10^{-3}$
(b) $c=10^{-5}$
We observe that strong regularization parameter $c=10^{-3}$ leads to the fast convergence of the $\mathcal{H}^{2}$-MG method for both matrices with $\sigma=0.1$ and $\sigma=0.01$, while the weak regularization parameter $c=10^{-5}$ leads to the divergence of both algorithms for the larger problem sizes.
@@ 4.2 Exponential kernel
4.2.1 Exponent matrix, asymptotics analysis
@@
We can see that strong regularization parameter $c=10^{-3}$ leads to the fast convergence of the $\mathcal{H}^{2}$-MG method for both matrices with $\sigma=0.1$ and $\sigma=0.01$, while the weak regularization parameter $c=10^{-5}$ leads to the divergence of both algorithms for the larger problem sizes.
@@
Figures 12(a) and 15 show the comparison of the overall solution time for the $\mathcal{H}^{2}$-MG and CG methods, compared against the H2-direct solver FMM-LU. A red cross indicates that either the iterative method failed to converge within 5000 iterations, or the direct solver failed to solve the system with the required accuracy.
@@ 5 Conclusion
The $\mathcal{H}^{2}$-MG algorithm offers an advance in solving large, dense kernel matrices efficiently by iterative means. By combining the rapid convergence of the multigrid method with the time and memory efficiencies of $\mathcal{H}^{2}$ matrix approximations, this algorithm not only fills a gap in the existing suite of $\mathcal{H}^{2}$ solvers but also expands the toolkit available for tackling complex computational problems. The demonstrated linear complexity and practical effectiveness of $\mathcal{H}^{2}$-MG , verified through numerical examples, underscore its potential as a for applications burdened by large, dense, kernel matrices. Future work will aim to expand its applicability beyond symmetric positive definite matrices, high-performance implementation, and integration into other computational frameworks.

### v2

@@ 3.4 Complexity analysis
The crucial feature of the $\mathcal{H}^{2}$-MG complexity analysis is the linear complexity of the $\mathcal{H}^{2}$ matrix. According to Remark 1, for the $\mathcal{H}^{2}$ matrix of size $N\times N$, the construction complexity, the memory requirements, and matrix-by-vector product complexity are all $\mathcal{O}(N)$. Assume $c_{H2}$ to be the $\mathcal{H}^{2}$ matrix-vector product constant.
We first compute the $\mathcal{H}^{2}$-MG computational complexity $n_{\text{op}}$ to run one V-cycle. The complexity of the fine grid smoothing iterations is $n_{f}c_{H2}N$, the coarse grid smoothing iterations is $n_{c}c_{H2}N_{i}$, for $i=2,\dots,l-1$, the complexity of the direct solver is $c_{d}N_{l}^{3}$, where $c_{d}$ is the direct solver complexity constant, and the restriction and prolongation operator complexity is negligible, compared to the smoothing iterations. The overall complexity is:
@@
Assume, for simplicity, that the block size is $B$ for all blocks and the block rank $r$ is fixed for all levels (consider it to be the maximum rank of any block). Also, assume the number of blocks of the initial matrix $A$ is $M$, and the number of blocks stacked together while transferring from level to level is $d$. Thus, $N=MB$, $N_{i}=\frac{Mr}{d^{i-2}}$, for $i=2,\dots,l$, and the overall complexity is:
@@
Using the sum of a geometric progression, we obtain:
@@
We next compute the memory requirements. The $\mathcal{H}^{2}$-MG algorithm requires storage of matrices $A$, $A_{i}$ for $i=2,\dots,l$, $U_{i}$, and $V_{i}$ for $i=1,\dots,l-1$. Matrices $U_{i}$ and $V_{i}$ are block-diagonal, and their storage is negligible compared to matrices $A$ and $A_{i}$. Assume $\mathcal{H}^{2}$ memory constant to be $c_{m}$. Thus, the memory requirements are:
@@
Analogously to the time complexity, we rewrite matrix sizes in terms of $M$, $B$, $r$, and $d$ and sum the geometric progression to obtain:
@@
Taking into account that $N_{l}$ is a constant, we obtain:
@@ 4 Numerical results
The numerical experiments showcase our Python implementation of the $\mathcal{H}^{2}$-MG algorithm, demonstrating $\mathcal{H}^{2}$-MG convergence for two different kernels and a boundary element method (BEM) example. The matrices are cast into the $\mathcal{H}^{2}$ format using the MCBH [30, 29] method. The code for the $\mathcal{H}^{2}$-MG algorithm is publicly available at https://github.com/dsushnikova/h2mg.
We compare the $\mathcal{H}^{2}$-MG algorithm with solvers CG and FMM-LU [37]. All solvers are implemented in Python without parallelism, and the experiments are run on a MacBook Pro (Apple M1 Pro, 16 GB RAM). All three methods benefit from the $\mathcal{H}^{2}$ structure of matrix $A$; thus, the comparison is natural. The $\mathcal{H}^{2}$-MG algorithm uses CG iteration for smoothing, illustrating how the coarse-level iterations improve the convergence relative to CG alone.
In the kernel matrices tests, we randomly generate $x_{\text{true}}$, then compute $b=Ax_{\text{true}}$ and solve the system $Ax=b$ for $x$. Then we plot convergence of the $A$-norm of the error $||e_{k}||_{A}$ vs. iteration count $k$, where the $A$-norm refers to the energy norm induced by matrix $A$.
@@
The right-hand sides were chosen in this manner so that the error norm at each iteration could be easily computed. We also tested a few cases where the right-hand sides were chosen randomly from a standard Gaussian distribution. In these cases, the linear systems were harder to solve, but $\mathcal{H}^{2}$-MG outperformed unpreconditioned CG, just as in the detailed results that we show below.
@@ 4.1.1 Gaussian matrix, analysis of error evolution
In our tests, we randomly generate $x_{\text{true}}$, then compute $b=Ax_{\text{true}}$ and solve the system $Ax=b$ for $x$. During the process, we can compute the error $e_{1*}=x_{k}-x_{\text{true}}$ since we know $x_{\text{true}}$. In the classic multigrid algorithm, the error $e_{1*}$ should behave in the following way: for the fine grid iterations, the high-frequency components of $e_{1*}$ should decrease rapidly, and during the coarse grid iterations, successive lower frequency components should decrease in turn.
In this section, we study the behavior of the error $e_{1*}$ in the $\mathcal{H}^{2}$-MG method. Consider $U_{1}$, the transfer matrix from the finest level to the coarser one. $U_{1}$ has orthonormal columns. Assume that the matrix $Q_{1}$ has columns that span the complementary space.
To analyze the components of the error after smoothing in the basis of the interpolation matrix $U_{1}$, or in other words, the part of the error that will be projected to the coarser levels, we compute $\text{pinv}(U_{1})e_{1*}$. Similarly, for the part of the after-smoothing error in the basis of the matrix $Q_{1}$, $\text{pinv}(Q_{1})e_{1e}$.
We study the 1D Gaussian matrix with points equally spaced between 00 and $1$, $N=1024$, $\sigma=0.01$, $c=10^{-3}$. The results of this analysis are presented in Figure 5. The blue curves represent the error components after the first smoothing step in the initial $V$-cycle, the red curves are for the second V-cycle.
@@
Figure 5: Convergence comparison for two successive coarse iterations for the $\mathcal{H}^{2}$-MG method
The graphs clearly shows that the error component $\text{pinv}(U_{1})e_{1*}$ is larger than that in the error component $\text{pinv}(Q_{1})e_{1*}$ components, demonstrating that the CG smoothing can complement coarse grid correction effectively.
4.1.2 Gaussian kernel matrix, analysis of number of smoothing steps
@@
We first study the convergence rate of the $\mathcal{H}^{2}$-MG depending on the number of fine iterations. Figure 6 presents the convergence comparison for different numbers of fine iterations, $n_{f}$ to the accuracy $\varepsilon=10^{-9}$, where $\varepsilon=\frac{(eAe^{\top})^{\frac{1}{2}}}{||b||_{2}}<10^{-9}$. The number of coarse iterations is fixed $n_{c}=40$. Fine and coarse iteration comparisons are given for matrix size $N=8\times 10^{4}$. For $\mathcal{H}^{2}$-MG , we can track only the $A$-norm of the error on the finest level, which makes it especially useful for comparison with CG. In all figures below, we plot this error per finest level iterations (outer iterations). Since coarse-level iterations are not shown directly in these plots, we also include error-per-time plots to reflect the computational effort spent on coarser levels.
(a) $A$-norm of the error per coarse grid iteration
(b) $A$-norm of the error per time
@@
We can see that in terms of convergence per iteration, fewer fine grid iterations lead to better convergence. Figure 7 presents the convergence comparison to the accuracy $\varepsilon=10^{-9}$ for different numbers of coarse iterations, $n_{c}$. The number of fine iterations is fixed at $n_{f}=1$.
(a) $A$-norm of the error per coarse grid iteration
(b) $A$-norm of the error per time
@@
4.1.3 Gaussian kernel matrix, asymptotics analysis
In this subsection we consider the asymptotic behavior of the system (9) with matrix $A$ given by (10) for various combinations of parameters $c$ and $\sigma$. Figures 8 and 9 show the convergence of the $\mathcal{H}^{2}$-MG algorithm compared to CG across various problem sizes. We use a tolerance $\varepsilon=10^{-9}$ and the parameters $n_{f}=1$, $n_{c}=40$.
We observe that the strong regularization parameter $c=10^{-3}$ leads to the fast convergence of the $\mathcal{H}^{2}$-MG method for both matrices with $\sigma=0.1$ and $\sigma=0.01$, while the weak regularization parameter $c=10^{-5}$ leads to the divergence of both algorithms for the larger problem sizes.
@@ 4.1.3 Gaussian kernel matrix, asymptotics analysis
(a) $c=10^{-3}$
(b) $c=10^{-5}$
(a) $c=10^{-3}$
(b) $c=10^{-5}$
As in the previous example, the number of V-cycles required for convergence to the fixed accuracy does not grow significantly.
@@ 4.2 Exponential kernel
4.2.1 Exponential kernel matrix, asymptotics analysis
@@
We can see that the strong regularization parameter $c=10^{-3}$ leads to the fast convergence of the $\mathcal{H}^{2}$-MG method for both matrices with $\sigma=0.1$ and $\sigma=0.01$, while the weak regularization parameter $c=10^{-5}$ leads to the divergence of both algorithms for the larger problem sizes.
@@
Figures 14 and 15 show the comparison of the overall solution time for the $\mathcal{H}^{2}$-MG and CG methods, compared against the H2-direct solver FMM-LU. A red cross in Figure 14 indicates that either the iterative method failed to converge within 5000 iterations, or the direct solver failed to solve the system with the required accuracy.
@@
4.3 Boundary element method for a 3D electrostatic problem
For the third example, we consider Laplace’s equation in its integral form, solved using the Boundary Element Method (BEM) on a complex 3D surface. This formulation is widely used in electrostatics, capacitance computations, molecular solvation models (such as the Poisson–Boltzmann equation), and modeling interactions between charged surfaces.
Specifically, we consider the Laplace single-layer potential, which describes the potential $u(\mathbf{x})$ at a point $\mathbf{x}\in\Gamma\subset\mathbb{R}^{3}$, where $\Gamma$ is a given surface. The potential $u(\mathbf{x})$ arises from charges (or equivalent sources) $\sigma(\mathbf{y})$ on the same surface $\Gamma$. The Laplace single-layer potential formula is:
[…]
where $\|\mathbf{x}-\mathbf{y}\|$ is the Euclidean distance between source and observation points, and $dS_{\mathbf{y}}$ is the surface area element.
To solve this numerically, we discretize the surface $\Gamma$ into $N$ triangular elements with associated centroids $\{\mathbf{y}_{j}\}$, quadrature weights $w_{j}$, and evaluation points $\{\mathbf{x}_{i}\}$.
To account for the singularity of the kernel when $i=j$, we modify the diagonal entries using a geometrically motivated regularization. We define $R_{i}$ as the average distance from the centroid of triangle $i$ to its three vertices. The resulting matrix $A\in\mathbb{R}^{N\times N}$ represents the discretized integral operator, with entries defined as:
(12) $A_{ij}=\begin{cases}\displaystyle\frac{w_{j}}{4\pi\|\mathbf{x}_{i}-\mathbf{y}_{j}\|},&\text{if }i\neq j,\\[10.0pt] \displaystyle\frac{w_{i}}{4\pi R_{i}},&\text{if }i=j,\end{cases}$
where
$R_{i}=\frac{1}{3}\sum_{k=1}^{3}\|\mathbf{y}_{i}-\mathbf{v}_{ik}\|,$
and $\{\mathbf{v}_{i1},\mathbf{v}_{i2},\mathbf{v}_{i3}\}$ are the vertices of triangle $i$.
For our numerical experiment, we chose a complex 3D geometry: a curved torus shown in Figure 18. To construct the right-hand side of the system, we consider a point source located at $\mathbf{x}_{0}\in\mathbb{R}^{3}$ outside the surface $\Gamma$. The potential generated by this point source at a location $\mathbf{y}\in\Gamma$ is given by the free-space Green’s function:
$f(\mathbf{y})=\frac{1}{4\pi\|\mathbf{y}-\mathbf{x}_{0}\|}.$
We evaluate this expression at each centroid $\mathbf{y}_{j}$ of the surface elements to obtain the right-hand side vector $\mathbf{f}\in\mathbb{R}^{N}$:
$f_{j}=\frac{1}{4\pi\|\mathbf{y}_{j}-\mathbf{x}_{0}\|},\quad j=1,\dots,N.$
We compare the CG method with $\mathcal{H}^{2}$-MG method on this problem. The comparison is performed for several mesh resolutions on the curved torus geometry. We set the number of fine grid iterations to $n_{f}=1$ and the number of coarse grid iterations to $n_{c}=40$. Figures 16(a) and 16(b) show the convergence behavior of both methods in terms of the number of iterations and total computational time, respectively. In this example, we plot the residual, since the exact solution is unknown and we cannot compute the $A$-norm of the error, as we did in the previous examples.
(a) Residual per fine-level iteration
(b) Residual per time
Figure 16: Convergence evolution of $\mathcal{H}^{2}$-MG and CG as problem size increases
We can see from the Figure 16(a) that the number of iterations required to solve the system using $\mathcal{H}^{2}$-MG does not increase significantly with problem size, while for CG it grows substantially. To confirm this effect quantitatively, Table 3 reports the number of V-cycles for $\mathcal{H}^{2}$-MG and the number of iterations for CG. Note that these numbers should not be compared directly, as one V-cycle includes several inner CG iterations. Instead, we focus on the growth trend.
Method Problem Size
6e3 12e3 26e3 50e3 100e3
$\mathcal{H}^{2}$-MG 1 2 2 2 3
CG 95 155 255 380 590
Table 3: Comparison of V-Cycles of $\mathcal{H}^{2}$-MG and iterations of CG across problem sizes
Figure 17 shows the total solution time for the system using CG and $\mathcal{H}^{2}$-MG . We observe that $\mathcal{H}^{2}$-MG outperforms CG in timing not only in terms of the constant factor but also in asymptotic scaling with problem size. Figure 18 presents the computed surface charge density $\sigma$ on the curved torus $\Gamma$. This example demonstrates that the $\mathcal{H}^{2}$-MG algorithm can be effectively applied to BEM problems.
Figure 17: Total solution time for CG and $\mathcal{H}^{2}$-MG across mesh sizes
Figure 18: Surface charge $\sigma$ distribution on the triangulated surface $\Gamma$ of a curved torus
@@ 5 Conclusion
The $\mathcal{H}^{2}$-MG algorithm offers an advance in solving large, dense kernel matrices efficiently by iterative means. By combining the rapid convergence of the multigrid method with the time and memory efficiencies of $\mathcal{H}^{2}$ matrix approximations, this algorithm not only fills a gap in the existing suite of $\mathcal{H}^{2}$ solvers but also expands the toolkit available for tackling complex computational problems. The demonstrated linear complexity and practical effectiveness of $\mathcal{H}^{2}$-MG, verified through numerical examples, underscores its potential for applications burdened by large, dense, kernel matrices. Future work will aim to expand its applicability beyond symmetric positive definite matrices, high-performance implementation, and integration into other computational frameworks.